bj can overhaul a boat's engine in 30 hours. his apprentice takes 60 hours to do the same job. how long would it take them to work together assuming no gain or loss in effeincy

Answer by richwmiller(17219)

1/a+1/b=1/x
1/30+1/60=1/x
60/1800+30/1800=1/x
90/1800=1/x
x=1800/90
x=20 hours working together
check
x/a+x/b=1
x/30+x/60=1
20/30+20/60=1
0.66666667+0.33333333=1
1=1
ok
codewt